OpenStudy (anonymous):

Using complete sentences, explain each step in simplifying the ratio x^2-x-20 __________ x^2+8x+16 ane simplified form. @Mathematics

OpenStudy (anonymous):

x^2-x-20 = (x-5)(x+4) x^2+8x+16=(x+4)(x+4) (x-5)(x+4)/(x+4)(x+4)=(x-5)/(x+4)